# Infinite-Dimensional Stochastic Differential Equations and Diffusion Dynamics of Coulomb Random Point Fields

### v3

@@ A.2 $\sum_{k=0}^{\infty}k\mu(S_{R}^{k})<\infty$
We set $\overline{S}^{m}_{\mathsf{r}}=\{{x}\in\mathbb{R}^{d};|{x}|\leq\mathsf{r}\}^{m}$. Let $j,k,l=1,\ldots,m$ and
$S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})=\big{\{}\mathbf{x}\in S_{\mathsf{r}}^{m}\,;\,\inf_{j\not=k}|{x}^{j}-{x}^{k}|>2^{-\mathsf{p}},\inf_{l,i}|{x}^{l}-s^{i}|>2^{-\mathsf{p}}\big{\}},$
$\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s})=\big{\{}\mathbf{x}\in\overline{S}^{m}_{\mathsf{r}}\,;\,\min_{j\not=k}|{x}^{j}-{x}^{k}|\geq 2^{-\mathsf{p}},\ \inf_{l,i}|{x}^{l}-s^{i}|\geq 2^{-\mathsf{p}}\big{\}}.$ (13.19)
Then, $S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$ is an open set and $\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s})$ is its closure in $(\mathbb{R}^{d})^{m}$.
Let $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ be a sequence of increasing sequences introduced before (13.16). For $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})$, we set
$\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}^{\circ}=\big{\{}(\mathbf{x},\mathsf{s})\in\mathsf{S}_{\mathrm{s,i}}^{[m]}\,;\,\ \mathbf{x}\in S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s}),\,\mathsf{s}\in\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}]\big{\}},$ (13.20)
$\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}=\big{\{}(\mathbf{x},\mathsf{s})\in\mathsf{S}_{\mathrm{s,i}}^{[m]}\,;\,\ \mathbf{x}\in\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s}),\ \mathsf{s}\in\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}]\big{\}}.$ (13.21)
Then, $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$ is an open set and $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$ is the closure of $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$. We easily see that $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$ is a compact set. Although $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$ and $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$ depend on $m\in\mathbb{N}$, we omit $m$ from the notation. Let
$\mathfrak{H}[\mathbf{a}]=\bigcup_{\mathsf{m}\in\mathbb{N}^{3}}\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}=\bigcup_{\mathsf{m}\in\mathbb{N}^{3}}\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}.$ (13.22)
We set $\mathbf{N}=\mathbb{N}\cup\mathbb{N}^{2}\cup\mathbb{N}^{3}$. For $\mathsf{m}\in\mathbf{N}$, we define $\mathsf{m}+1\in\mathbf{N}$ as
$\mathsf{m}+1=\begin{cases}(\mathsf{p}+1,\mathsf{q},\mathsf{r})&\text{ for $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}$, }\\ (\mathsf{q}+1,\mathsf{r})&\text{ for $\mathsf{m}=(\mathsf{q},\mathsf{r})\in\mathbb{N}^{2}$, }\\ \mathsf{r}+1&\text{ for $\mathsf{m}=\mathsf{r}\in\mathbb{N}$.}\end{cases}$ (13.23)
We set $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}=\cup_{\mathsf{p}=1}^{\infty}\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ for $\mathsf{m}=(\mathsf{q},\mathsf{r})\in\mathbb{N}^{2}$ and $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}=\cup_{\mathsf{q}=1}^{\infty}\cup_{\mathsf{p}=1}^{\infty}\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ for $\mathsf{m}=(\mathsf{r})\in\mathbb{N}$. We set $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$ analogously. Clearly, for all $\mathsf{m}\in\mathbf{N}$
$\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}\subset\mathfrak{H}[\mathbf{a}]_{\mathsf{m}},\quad\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}\subset\mathfrak{H}[\mathbf{a}]_{\mathsf{m}+1}.$
Cut-off functions $\chi_{\mathsf{m}}$ on $(\mathbb{R}^{d})^{m}\times\mathsf{S}$: Let $\tilde{\varphi}_{\mathsf{r}}\in C_{0}^{\infty}(\mathbb{R})$ be such that
$\tilde{\varphi}_{\mathsf{r}}(t)=\begin{cases}1&\text{for $|t|\leq\mathsf{r}$}\\ 0&\text{for $|t|\geq\mathsf{r}+1$},\end{cases}\quad 0\leq\tilde{\varphi}_{\mathsf{r}}(t)\leq 1,\quad\lvert\tilde{\varphi}_{\mathsf{r}}^{\prime}(t)\rvert\leq 2\quad\text{ for all $t\in\mathbb{R}$}.$
@@
$h_{\mathsf{p}}(t)=\begin{cases}1&(t\leq 2^{-\mathsf{p}-2}),\\ 0&(2^{-\mathsf{p}-1}\leq t)\end{cases},\quad 0\leq h_{\mathsf{p}}(t)\leq 1,\quad|h_{\mathsf{p}}^{\prime}(t)|\leq 2^{\mathsf{p}+3}\quad\text{ for all }t.$
We write $\mathbf{x}=({x}^{k})_{k=1}^{m}\in(\mathbb{R}^{d})^{m}$ and $\mathsf{s}=\sum_{i}\delta_{s^{i}}\in\mathsf{S}$. Let $h_{\mathsf{p}}^{[m]}\!:\!(\mathbb{R}^{d})^{m}\times\mathsf{S}\!\to\!\mathbb{R}$ such that
$\quad\quad\quad h_{\mathsf{p}}^{[m]}(\mathbf{x},\mathsf{s})=\prod_{k=1}^{m}\Big{\{}\prod_{j\not=k}^{m}\{1-h_{\mathsf{p}}(|{x}^{k}-{x}^{j}|)\}\Big{\}}\Big{\{}\prod_{i}\{1-h_{\mathsf{p}}(|{x}^{k}-s^{i}|)\}\Big{\}}.$
We label $\mathsf{s}=\sum_{i}\delta_{s^{i}}$ in such a way that $|s^{i}|\leq|s^{i+1}|$ for all $i$. Let $a_{\mathsf{q}}=\{a_{\mathsf{q}}(\mathsf{r})\}_{\mathsf{r}\in\mathbb{N}}$ be the increasing sequences in (13.16). We set for $Q\in\mathbb{N}\cup\{\infty\}$
$\mathit{d}_{\mathsf{q}}^{Q}(\mathsf{s})=\Big{\{}\sum_{R=1}^{Q}\sum_{s^{i}\in\overline{S}_{R},i>a_{\mathsf{q}}(R)}(R-|s^{i}|)^{2}\Big{\}}^{1/2}.$ (13.24)
We introduce the cut-off functions defined as
$\chi_{\mathsf{r}}(\mathbf{x},\mathsf{s})$ $=\varphi_{\mathsf{r}}(\mathbf{x}),$
$\chi_{\mathsf{q},\mathsf{r}}(\mathbf{x},\mathsf{s})$ $=\varphi_{\mathsf{r}}(\mathbf{x})(h_{0}\circ\mathit{d}_{\mathsf{q}}^{\infty})(\mathsf{s}),$
$\chi_{\mathsf{p},\mathsf{q},\mathsf{r}}(\mathbf{x},\mathsf{s})$ $=\varphi_{\mathsf{r}}(\mathbf{x})(h_{0}\circ\mathit{d}_{\mathsf{q}}^{\infty})(\mathsf{s})h_{\mathsf{p}}^{[m]}(\mathbf{x},\mathsf{s}).$
We set $\chi_{\mathsf{m}}=\chi_{\mathsf{p},\mathsf{q},\mathsf{r}}$. Then, $\{\chi_{\mathsf{m}}\}$ are consistent in the sense that $\chi_{\mathsf{m}}(\mathbf{x},\mathsf{s})=\chi_{\mathsf{m}+1}(\mathbf{x},\mathsf{s})$ for $(\mathbf{x},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$. We suppress $m$ from the notation of $\chi_{\mathsf{m}}$ although $\chi_{\mathsf{m}}$ depends on $m\in\mathbb{N}$. We take the limit in $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})$ along with the order $\mathsf{m}\mapsto\mathsf{m}+1$ in (13.23). Let
$\lim_{\mathsf{m}}:=\lim_{\mathsf{r}\to\infty}\lim_{\mathsf{q}\to\infty}\lim_{\mathsf{p}\to\infty}.$ (13.25)
Then, we easily see that
$\lim_{\mathsf{m}}\chi_{\mathsf{m}}(\mathbf{x},\mathsf{s})=1\quad\text{ for all $(\mathbf{x},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]$}.$
By a direct calculation similar to that in [22, Lem. 2.5], we obtain the following.
Lemma 13.6.
For each $m\in\mathbb{N}$, the functions $\chi_{\mathsf{m}}$, $\mathsf{m}\in\mathbf{N}$, satisfy the following.
$\chi_{\mathsf{m}}(\mathbf{x},\mathsf{s})=\begin{cases}0&\text{ for }(\mathbf{x},\mathsf{s})\not\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}+1}\\ 1&\text{ for }(\mathbf{x},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}\end{cases},\quad\chi_{\mathsf{m}}\in C_{0}((\mathbb{R}^{d})^{m})\otimes\mathscr{D}_{\bullet\mathrm{b}}^{[1]},$
[…]
13.7. A sufficient condition for ((IFC)).
Let $(\mathbf{X},\mathbf{B})$ be a weak solution of (13.14) defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$. For $\mathbf{X}=(X^{i})_{i\in\mathbb{N}}$, we set the $m$-labeled process $\mathbf{X}^{[m]}=(\mathbf{X}^{m},\mathsf{X}^{m*})$ such that $\mathbf{X}^{m}=(X^{1},\ldots,X^{m})$ and $\mathsf{X}_{t}^{m*}=\sum_{j=m+1}^{\infty}\delta_{X_{t}^{j}}$. Let $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$ be as in (13.20). Let $\lim_{\mathsf{m}}$ be as in (13.25).
We make an assumption.
[…]
$P(\lim_{\mathsf{m}}\varsigma_{\mathsf{m}}(\mathbf{X}^{[m]})=\infty)=1,$
where $\varsigma_{\mathsf{m}}(\mathbf{w}^{[m]})=\inf\{t>0;\mathbf{w}^{[m]}_{t}\not\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}\}$ is the exit time from $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}^{\circ}$.
@@
Let $S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$ be as in (13.19). For $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}$ and $(\mathbf{x},\mathsf{s}),(\mathbf{y},\mathsf{s})\in S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$, we set $(\mathbf{x},\mathsf{s})\sim_{\mathsf{m}}(\mathbf{y},\mathsf{s})$ if $\mathbf{x}$ and $\mathbf{y}$ are in the same connected component of $S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$ and $\mathsf{s}\in\Pi_{2}(\mathfrak{H}[\mathbf{a}]_{\mathsf{m}})$. Here, $\Pi_{2}$ is a projection $\Pi_{2}\!:\!(\mathbb{R}^{d})^{m}\times\mathsf{S}\!\to\!\mathsf{S}$ given by $\Pi_{2}(\mathbf{x},\mathsf{s})=\mathsf{s}$ and $\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ is denoted by (13.21). We extend the domain of $\mathfrak{u}$ from $(\mathbb{R}^{d})^{m}$ to $(\mathbb{R}^{d})^{m}\times\mathsf{S}$ such that $\mathfrak{u}^{[m,0]}(\mathbf{x},\mathsf{s})=\mathfrak{u}(\mathbf{x})+\mathsf{s}$.
Let $\mathsf{S}_{\mathrm{sde}}$ be the set defined before (13.14). By definition, $\mathsf{S}_{\mathrm{sde}}$ is the set such that the coefficient of (13.14) is defined on $\mathsf{S}_{\mathrm{sde}}^{[1]}=(\mathfrak{u}^{[1,0]})^{-1}(\mathsf{S}_{\mathrm{sde}})$.
@@
ii For each $\mathsf{m}=(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}$, there exists a constant $\mathsf{F}_{\mathsf{m}}$ such that
$|\sigma^{m}(\mathbf{x},\mathsf{s})-\sigma^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{m}}|\mathbf{x}-\mathbf{y}|,$ (13.27)
$|\mathsf{b}^{m}(\mathbf{x},\mathsf{s})-\mathsf{b}^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{m}}|\mathbf{x}-\mathbf{y}|$ (13.28)
for all $(\mathbf{x},\mathsf{s}),(\mathbf{y},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]_{\mathsf{m}}$ such that $(\mathbf{x},\mathsf{s})\sim_{\mathsf{m}}(\mathbf{y},\mathsf{s})$.
@@
Suppose that (13.14) has a weak solution $(\mathbf{X},\mathbf{B})$ satisfying $\{\mathbf{B1}\}$ and that $\{\mathbf{B2}\}$ holds. Then, $(\mathbf{X},\mathbf{B})$ satisfies ((IFC)).
@@
[…]
Let $\nu$ be a random point field such that $\nu(\mathsf{S}_{\mathrm{sde}})=1$ and let $\mathfrak{l}\!:\!\mathsf{S}_{\mathrm{s,i}}\!\to\!(\mathbb{R}^{d})^{\mathbb{N}}$ be a label. We shall consider ISDE (13.14) with the initial distribution $\nu\circ\mathfrak{l}^{-1}$. Let $\{Q_{\mathsf{s}}\}$ be a family of probability measures on $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\})$ such that $(\mathbf{X},\mathbf{B})$ defined on $(\Omega,\mathcal{F},Q_{\mathsf{s}},\{\mathcal{F}_{t}\})$ is a weak solution of (13.14) starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$ for $\nu$-a.s. $\mathsf{s}$. We assume $\{Q_{\mathsf{s}}\}$ is a measurable family in the following sense.
[…]
Let $Q_{\mathbf{x},\mathsf{s}}^{[m]}$ be the distribution of $\mathbf{X}^{[m]}=(\mathbf{X}^{m},\mathsf{X}^{m*})$ under $Q_{\mathfrak{u}(\mathbf{x})+\mathsf{s}}$. Then,
$\int_{C([0,\infty);(\mathbb{R}^{d})^{m}\times\mathsf{S})}f(\mathbf{w}_{t}^{[m]})dQ_{\mathbf{x},\mathsf{s}}^{[m]}=\int f(\mathbf{X}^{[m]}_{t})dQ_{\mathfrak{u}(\mathbf{x})+\mathsf{s}}.$
Let $\nu^{[m]}$ be the $m$-reduced Campbell measure of $\nu$ for $m\in\mathbb{N}$. Let $\nu^{[0]}=\nu$. We set
$Q_{\nu}^{[m]}=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}Q_{\mathbf{x},\mathsf{s}}^{[m]}d\nu^{[m]}\quad\text{ for $m\in\{0\}\cup\mathbb{N}$}.$
By definition, $Q_{\nu}^{[0]}=Q_{\nu}\circ\mathsf{X}^{-1}$. We make an assumption.
@@
Let $\theta\in C^{\infty}(\mathbb{R})$ such that $0\leq\theta(t)\leq 1$ for all $t\in\mathbb{R}$, $\theta(t)=0$ for $t\leq\epsilon$, and $\theta(t)=1$ for $t\geq 1-\epsilon$ for a sufficiently small $\epsilon>0$. Assume that $0\leq\theta^{\prime}(t)\leq\sqrt{2}$ for all $t\in\mathbb{R}$. Let $\mathit{d}_{\mathsf{q}}^{Q}$ be the function on $\mathsf{S}$ defined by (13.24). Let
@@
The limit in (13.29) exists because $\theta\circ\mathit{d}_{\mathsf{q}}^{Q}(\mathsf{s})$ is non-negative and non-decreasing in $Q$. We make an assumption.
@@
In practice, (13.30) is easy to check. Indeed, we have the following.
Lemma 13.7 ([16, Lem. 5.4]).
(13.30) follows from (13.17) and
$\sum_{\mathsf{q}=1}^{\infty}\mathsf{q}^{2}\nu(\mathsf{K}[a_{\mathsf{q}}]^{c})<\infty.$ (13.31)
Lemma 13.8 ([16, Prop. 6.3]).
Assume that $\{\mathbf{UB}\}$, $\{\mathbf{MF}\}$, $\{\mathbf{BX}\}$, $\{\mathbf{S}_{\nu}\}$, and (13.30) hold. Let $\kappa_{\mathsf{q}}=\inf\{t>0;\mathsf{X}_{t}\notin\mathsf{K}[a_{\mathsf{q}}]\}$ Then, $P(\lim_{\mathsf{q}\to\infty}\kappa_{\mathsf{q}}=\infty)=1$.

### v4

@@ 13.7. Cut-off compact sets and a sufficient condition for ((IFC))
We set $\overline{S}^{m}_{\mathsf{r}}=\{{x}\in\mathbb{R}^{d};|{x}|\leq\mathsf{r}\}^{m}$. Let $j,k,l=1,\ldots,m$. Let
$\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s})=\big\{\mathbf{x}\in\overline{S}^{m}_{\mathsf{r}}\,;\,\min_{j\not=k}|{x}^{j}-{x}^{k}|\geq 2^{-\mathsf{p}},\ \inf_{l,i}|{x}^{l}-s^{i}|\geq 2^{-\mathsf{p}}\big\},$
$S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})=\big\{\mathbf{x}\in S_{\mathsf{r}}^{m}\,;\,\inf_{j\not=k}|{x}^{j}-{x}^{k}|>2^{-\mathsf{p}},\inf_{l,i}|{x}^{l}-s^{i}|>2^{-\mathsf{p}}\big\}.$ (13.10)
Then $S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$ is an open set and $\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s})$ is its closure in $(\mathbb{R}^{d})^{m}$.
Let $\mathbf{a}=\{a_{\mathsf{q}}\}_{\mathsf{q}\in\mathbb{N}}$ and $\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}]$ be as in (13.3)–(13.5). For $\mathsf{k}=(\mathsf{p},\mathsf{q},\mathsf{r})$, let
$\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ $=\big\{(\mathbf{x},\mathsf{s})\in\mathsf{S}_{\mathrm{s,i}}^{[m]}\,;\,\ \mathbf{x}\in\overline{S}^{m}_{\mathsf{p},\mathsf{r}}(\mathsf{s}),\ \mathsf{s}\in\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}]\big\},$
$\mathfrak{H}[\mathbf{a}]$ $=\bigcup_{\mathsf{p},\mathsf{q},\mathsf{r}\in\mathbb{N}}\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}.$ (13.11)
Although $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}=\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ depends on $m\in\mathbb{N}$, we suppress the dependence on $m$ in the notation. Clearly, the set $\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$ is increasing in each parameter $\mathsf{p},\mathsf{q},\mathsf{r}\in\mathbb{N}$. Hence, by monotonicity, the limit of the following exit times $\varsigma_{\mathsf{p},\mathsf{q},\mathsf{r}}$ as $\mathsf{p},\mathsf{q},\mathsf{r}\to\infty$, taken successively in $\mathsf{p}$, $\mathsf{q}$, and $\mathsf{r}$, is well defined.
Let $\mathsf{S}_{\mathrm{sde}}$ be the set defined before (13.7). By definition, $\mathsf{S}_{\mathrm{sde}}$ is the set such that the coefficient of (13.7) is defined on $\mathsf{S}_{\mathrm{sde}}^{[1]}=(\mathfrak{u}^{[1,0]})^{-1}(\mathsf{S}_{\mathrm{sde}})$. Here we extend the domain of $\mathfrak{u}$ to $(\mathbb{R}^{d})^{m}\times\mathsf{S}$ such that $\mathfrak{u}^{[m,0]}(\mathbf{x},\mathsf{s})=\mathfrak{u}(\mathbf{x})+\mathsf{s}$.
Let $(\mathbf{X},\mathbf{B})$ be a weak solution of (13.7) defined on $(\Omega,\mathscr{F},P,\{\mathscr{F}_{t}\})$. Let $\mathbf{X}^{[m]}$ be the associated $m$-labeled process. Let $\nu$ be the RPF in ((AC))ν.
@@
i $\nu(\mathsf{S}\backslash\bigcup_{\mathsf{q}}\mathsf{K}[\mathit{a}_{\mathsf{q}}^{+}])=0$.
ii The inclusion $\mathfrak{u}^{[m,0]}(\mathfrak{H}[\mathbf{a}])\subset\mathsf{S}_{\mathrm{sde}}$ holds.
iii $\mathbf{X}^{[m]}=(\mathbf{X}^{m},\mathsf{X}^{m*})$ does not exit from $\mathfrak{H}[\mathbf{a}]=\bigcup_{(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}}\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$:
$P(\lim_{\mathsf{r}\to\infty}\lim_{\mathsf{q}\to\infty}\lim_{\mathsf{p}\to\infty}\varsigma_{\mathsf{p},\mathsf{q},\mathsf{r}}(\mathbf{X}^{[m]})=\infty)=1.$
Here $\varsigma_{\mathsf{p},\mathsf{q},\mathsf{r}}(\mathbf{w}^{[m]})=\inf\{t>0;\mathbf{w}^{[m]}_{t}\not\in\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}\}$ is the exit time from $\mathfrak{H}[\mathbf{a}]_{\mathsf{p},\mathsf{q},\mathsf{r}}$.
@@
For $\mathsf{k}=(\mathsf{p},\mathsf{q},\mathsf{r})\in\mathbb{N}^{3}$ and $(\mathbf{x},\mathsf{s}),(\mathbf{y},\mathsf{s})\in S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$, we set $(\mathbf{x},\mathsf{s})\sim_{\mathsf{k}}(\mathbf{y},\mathsf{s})$ if $\mathbf{x}$ and $\mathbf{y}$ are in the same connected component of $S_{\mathsf{p},\mathsf{r}}^{m}(\mathsf{s})$ and $\mathsf{s}\in\Pi_{2}(\mathfrak{H}[\mathbf{a}]_{\mathsf{k}})$. Here, $\Pi_{2}$ is a projection $\Pi_{2}\!:\!(\mathbb{R}^{d})^{m}\times\mathsf{S}\!\to\!\mathsf{S}$ given by $\Pi_{2}(\mathbf{x},\mathsf{s})=\mathsf{s}$.
@@
$|\sigma^{m}(\mathbf{x},\mathsf{s})-\sigma^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{k}}|\mathbf{x}-\mathbf{y}|,$ (13.13)
$|\mathfrak{b}^{m}(\mathbf{x},\mathsf{s})-\mathfrak{b}^{m}(\mathbf{y},\mathsf{s})|\leq\mathsf{F}_{\mathsf{k}}|\mathbf{x}-\mathbf{y}|$ (13.14)
for all $(\mathbf{x},\mathsf{s}),(\mathbf{y},\mathsf{s})\in\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$ such that $(\mathbf{x},\mathsf{s})\sim_{\mathsf{k}}(\mathbf{y},\mathsf{s})$, where $\mathfrak{H}[\mathbf{a}]_{\mathsf{k}}$ is the set defined in (13.11) and $\sim_{\mathsf{k}}$ is the equivalence relation defined after (13.12).
@@
Suppose that (13.7) has a weak solution $(\mathbf{X},\mathbf{B})$ satisfying $\{\mathbf{B1}\}$ and that $\{\mathbf{B2}\}$ holds. Then $(\mathbf{X},\mathbf{B})$ satisfies ((IFC)).
@@
Theorem 13.5 is essentially proved in [34, Prop. 11.1]. Since the formulation there is slightly different, we explain the correspondence between the assumptions.
[…]
In [34, Prop. 11.1], it is further assumed that the coefficient $\sigma^{m}$ of the ISDE is constant. This assumption can, however, be removed. Indeed, it suffices to use [34, Lem. 11.2]1 in place of [34, Lem. 11.2]3 in the proof, which can be verified directly. ∎
[…]
Let $a=\sigma{}^{t}\sigma$, where $\sigma$ is as in (13.7).
[…]
Let $\nu$ be an RPF such that $\nu(\mathsf{S}_{\mathrm{sde}})=1$. Let $\nu^{[m]}$ be the $m$-reduced Campbell measure of $\nu$ for $m\in\mathbb{N}$ and $\nu^{[0]}=\nu$.
Let $\mathfrak{l}\!:\!\mathsf{S}_{\mathrm{s,i}}\!\to\!(\mathbb{R}^{d})^{\mathbb{N}}$ be a label. Let $\{Q_{\mathsf{s}}\}$ be a family of probability measures on $(\Omega,\mathcal{F},\{\mathcal{F}_{t}\})$ such that $(\mathbf{X},\mathbf{B})$ defined on $(\Omega,\mathcal{F},Q_{\mathsf{s}},\{\mathcal{F}_{t}\})$ is a weak solution of (13.7) starting at $\mathbf{s}=\mathfrak{l}(\mathsf{s})$ for $\nu$-a.s. $\mathsf{s}$.
We assume $\{Q_{\mathsf{s}}\}$ is a measurable family in the following sense.
@@
Let $Q_{\mathbf{x},\mathsf{s}}^{[m]}$ be the distribution of $\mathbf{X}^{[m]}=(\mathbf{X}^{m},\mathsf{X}^{m*})$ under $Q_{\mathfrak{u}(\mathbf{x})+\mathsf{s}}$. We define $Q_{\nu}^{[m]}=\int_{(\mathbb{R}^{d})^{m}\times\mathsf{S}}Q_{\mathbf{x},\mathsf{s}}^{[m]}d\nu^{[m]}$ for $m\in\mathbb{N}$ and $Q_{\nu}^{[0]}=Q_{\nu}\circ\mathsf{X}^{-1}$.
@@
We label $\mathsf{s}=\sum_{i}\delta_{s^{i}}$ in such a way that $|s^{i}|\leq|s^{i+1}|$ for all $i$. Let $a_{\mathsf{q}}=\{a_{\mathsf{q}}(\mathsf{r})\}_{\mathsf{r}\in\mathbb{N}}$ be the increasing sequences in (13.3).
For $Q\in\mathbb{N}\cup\{\infty\}$, define
$\mathit{d}_{\mathsf{q}}^{Q}(\mathsf{s})=\Big\{\sum_{R=1}^{Q}\sum_{s^{i}\in\overline{S}_{R},i>a_{\mathsf{q}}(R)}(R-|s^{i}|)^{2}\Big\}^{1/2}.$
Let $\theta\in C^{\infty}(\mathbb{R})$ such that $0\leq\theta(t)\leq 1$ for all $t\in\mathbb{R}$, $\theta(t)=0$ for $t\leq\epsilon$, and $\theta(t)=1$ for $t\geq 1-\epsilon$ for a sufficiently small $\epsilon>0$. Assume that $0\leq\theta^{\prime}(t)\leq\sqrt{2}$ for all $t\in\mathbb{R}$. Let
@@
The limit in (13.15) exists since $\theta\circ\mathit{d}_{\mathsf{q}}^{Q}\geq 0$ is non-decreasing in $Q$.
Lemma 13.8 ([18, Lem. 5.4]).
Let $\mathsf{K}[a_{\mathsf{q}}]$ be as in (13.4). Assume (13.5) and $\sum_{\mathsf{q}=1}^{\infty}\mathsf{q}^{2}\nu(\mathsf{K}[a_{\mathsf{q}}]^{c})<\infty$. Then
@@
[…]
Lemma 13.9 ([18, Prop. 6.3]).
Assume $\{\mathbf{UB}\},\{\mathbf{MF}\},\{\mathbf{BX}\},\{\mathbf{S}_{\nu}\}$, and (13.16). Let $\mathsf{X}=\mathfrak{u}_{\mathrm{path}}(\mathbf{X})$ be the unlabeled dynamics under $Q_{\nu}$. Then
$P(\lim_{\mathsf{q}\to\infty}\kappa_{\mathsf{q}}=\infty)=1.$
Here $\kappa_{\mathsf{q}}=\inf\{t>0\,;\,\mathsf{X}_{t}\notin\mathsf{K}[a_{\mathsf{q}}]\}$.